Suppose that in a city there are 100 identical stores selling the same product. The total daily market demand function for the produce in the market is QD =250,000 – 10,000P, where P is expressed in dollars per unit. The daily market supply curve is QS = -50,000 + 20,000P

a. Determine algebraically the equilibrium price and quantity of product.

b. Now suppose the market is monopolized (e.g., a cartel is that determines price and quantity as a monopolist would). Determine the monopolist’s profit-maximizing price and quantity.

c. What is the dead weight loss (DWL)?
Answer #1

Equilibrium is achieved when Quantity demanded= quantity supplied

So 250,000-10,000P= -50000+20000P

250000+50000=20000P+10000P

750000=30000P

P=750000/30000

P=25